# Infra-red thermographic inversion in ST40

### v1

@@ Abstract
A new tool for infra-red thermographic inversion on Tokamak Energy’s spherical tokamak, ST40, is developed in-house and here presented. Functional Analysis of Heat Flux (FAHF) is written in Python, and configured for multi-2D thermographic inversions, solving the heat conduction equation within the divertor tiles via the finite difference method, and leveraging an explicit time-marching scheme. Using infra-red camera data of the highest effective spatial resolution available, FAHF ultimately outputs the plasma perpendicular heat flux on the divertor, crucial quantity in any edge plasma investigation. In the present work, the internal numerics of the tool is first successfully verified by formal time and space convergence analyses, and corroborated by an energy balance assessment. A significant sensitivity of the perpendicular heat flux computed by FAHF to the user-selected spatial resolution is then evidenced. However, a precise heat flux is proved to be recoverable by ensuring a sufficiently high resolution. Last, the appropriateness of the model/geometry simplifications adopted in FAHF is successfully confirmed, by means of comparison against COMSOL Multiphysics® simulations. FAHF is hence conclusively proven to qualify as a precise and accurate tool for infra-red thermographic inversions.
@@ Introduction
Tokamak-based nuclear fusion presents a plethora of physico-technological challenges to be overcome for its achievement [1]. A robust power exhaust (PEX) solution figures among them. This particular challenge arises from the substantial heat the fusion-grade core plasma leaks in the peripheral region of open magnetic field lines, the scrape-off layer (SOL). This power, expected to amount to hundreds of MW in a thermonuclear reactor [2, 3], is unusable by the plasma to generate further fusion reactions, and it is invariably channelled by the SOL out of the plasma system, towards the plasma-facing components (PFCs).
However, irrespective of the total available area offered by the PFCs, the peculiar anisotropicity of the SOL plasma transport [4, 5] favours a severely localised heat deposition [6], which nominally translates in heat flux densities in the order of few hundreds of $\text{MW}\cdot\text{m}^{-2}$ to be handled. Although mitigatory action is routinely undertaken in tokamaks [7, 8, 9], the baseline PFC technology currently available and tested has successfully demonstrated[10] suitable PEX properties and mechanical integrity only up to few tens of $\text{MW}\cdot\text{m}^{-2}$. Carefully determining the value attained by the parameters affecting the heat flux footprint hence becomes vital. Among them are the peak heat flux $q_{0}$ and power fall-off length $\lambda_{q}$, respectively dictating the severity of the flux, and the width of the footprint [6].
The well-established diagnostic method of infra-red (IR) thermography [11] offers the possibility of a detailed determination of such parameters, and of a thorough multi-dimensional characterisation of the SOL plasma at the interface with the PFCs. By providing a two-dimensional (2D) view of the thermally loaded PFCs, the evolution in time of the PFC surface temperature diagnosed by the IR camera is leveraged to infer the heat flux which caused such evolution to occur. This inference is known as thermographic inversion, and involves the solution of the heat conduction (Fourier) equation [12]. But, among others, the three-dimensionality of the PFCs [13], surface imperfections and deposited impurity layers[14, 15] significantly exacerbate the modelling complexity. A non-exhaustive list of teams reporting on IR thermographic apparati and/or thermographic inversion analyses which aids further insights is: Alcator C-Mod[16], ASDEX[17, 18, 19], COMPASS[20], DIII-D[21], EAST[22], ITER[23], JET[24, 25], KSTAR[26, 27], LHD[28], MAST[29, 30], NSTX[31], SPARC[32, 33], TCV[34, 35], W7-X[36], and WEST[37, 38, 39].
In the current work, the new in-house Python-based tool for the Functional Analysis of Heat Flux (FAHF) in the ST40 spherical tokamak [40, 41], owned and operated by Tokamak Energy Ltd., is presented. With the goal of constituting a fast and flexible inter-shot analysis tool, simplifications are accordingly adopted in FAHF. Therefore, in addition to the customary code verification to quantify the code’s precision, and assessment of the output physical parameters, the appropriateness of the simplifying assumptions of FAHF is scrutinised by means of comparison with the renowned COMSOL Multiphysics® modelling suite [42], to estimate its accuracy. Cross-code comparisons indeed constitute an effective method for the quantification of model and geometry errors [43, 44], paramount in the present context.
@@
The paper is organised as follows. Sec. 2 expounds on the ST40 tokamak, experimental setup and data selection. The numerical methods and tools adopted are described in Sec. 3. The results of the study are collated in Sec. 4, and discussed in Sec. 5. Lastly, Sec. 6 summarises the conclusion of the work, and comments on future studies.
@@
ST40 is a high-field compact spherical tokamak (ST) constructed and operated by Tokamak Energy Ltd., characterised by a major radius $R=0.4-0.5$ m, minor radius $a=0.2-0.25$ m, and a magnetic field up to 2.0 T on axis.
Disconnected double-null diverted plasmas are most commonly featured by ST40. Fig. 1a shows the EFIT magnetic reconstruction [45] in the poloidal cross section for shot 11376 at time instant 128 ms (Sec. 2.3).
ST40 is provided with highly-shaped fish-scaled[13] tiled divertor targets, pictured in Fig. 1b, with inter-tile gaps of 4.5 mm in width. The leading edge of each tile is poloidally chamfered[13], an area that is not of interest in the present study, and therefore neglected. The leftover volume of the divertor tiles is thus a rectangular cuboid, with 4 mm of plasma-facing TZM molybdenum (Mo) overlaid to 25 mm of copper-cromium-zirconium (CuCrZr). Poloidally, the geometry of the divertor is that of an open non-baffled divertor, with the inner leg orthogonal to the plate, and an horizontal outer leg (Fig. 1a).
The reader is referred to Ref. [41] for further information about ST40, and to Ref. [46] for specifics of its divertor.
@@
ST40 is equipped with a FLIR X6903sc MWIR ($3-5\;\mu$m) camera, which features an acquisition rate of 0.8 kHz ($\Delta t_{\text{IR}}=$ 1.25 ms) for its 612 $\times$ 540 pixels, and diagnoses the upper outer divertor target (Fig. 1a). The camera was installed in collaboration with Oak Ridge National Laboratory. Its line of sight (LoS) on the divertor provided by an ad-hoc re-entrant optical endoscope is close to perpendicular to the surface, and the field of view (FoV) covers the minimum amount of surface required to fully capture the heat flux footprint. The compenetration of the above LoS and FoV hence maximises the camera effective spatial resolution along the divertor, which results in a noteworthy $\Delta x_{\text{IR}}=0.21\pm 0.03$ mm/pixel ($\sim 5$ pixel/mm), appreciable in Fig. 1c. The comparison with the list of references in Sec. 1 actually suggests the spatial resolution of ST40 IR camera to be the highest available worldwide for thermographic inversions of a full plasma heat flux footprint.
The calibration of the IR camera and endoscope has been accomplished ex-situ via a known black-body emission source, in the temperature window $[20;118]$. The surface emissivity of the divertor tiles is instead assumed to be unitary, constant with respect to the position in space and surface temperature. This does not impact the main focus of the present work.
@@
Figure 1: (a) EFIT reconstruction of the disconnected double-null magnetic equilibrium of ST40’s shot 11376 at 128 ms: primary separatrix in magenta solid; secondary separatrix in magenta dashed; further magnetic surfaces in grey; wall in black; chord 10 in blue (1:1 scale). (b) IR camera view at 128 ms on the upper outer divertor in pixel space $(x_{\text{p}},\phi_{\text{p}})$. Overlaid are (see text for details): the FAHF reference frame in real space with $\phi$, proxy for toroidal direction, $x$, proxy for poloidal direction, and the direction in the depth of the tiles $y$ (entering the screen); an illustrative representation of chord 10 of the bundle in both its sub-chords (one per tile, blue dots). (c) Temperature $T_{\text{IR}}$ as extracted from the IR view via the Calcam software [47] (Appendix A) along chord 10 in real space, with the empty gap in its middle. The data-point resolution is the actual $0.21\pm 0.03$ mm/pixel of the ST40 IR camera.
@@
Shot number 11376 within the ST40 database is selected for the analyses of the present study. Reasons behind the choice are primarily the availability and quality of the IR experimental data.
Furthermore, shot 11376 is a representative instance of the ”average” ST40 plasma shot, from the edge plasma point of view. It is in fact a quiet L-mode, featuring a single-exponential heat flux footprint (Sec. 4.1.3).
@@
Additionally, during the time window $[50;140]$ ms the plasma presents a stable flat-top. Although the IR camera is marginally affected by saturation in $[130;140]$ ms in the region of interest (Sec. 3.3.1), this does not impact the numerical characterisation of FAHF.
@@
Despite the magnetic equilibrium of shot 11376 nominally being a lower-biased disconnected double-null (Fig. 1a), the low degree of disconnection ($\delta r_{\text{sep}}/\lambda_{q}\sim 1.5$ at the outer mid-plane [48]) allows the plasma to deposit a satisfactory amount of power in the camera FoV during the actual divertor heating phase in $[90;140]\;\text{ms}$.
@@
Theoretical fundamentals and model assumptions
Safely discarding any volumetric heat source/sink (e.g. arising from induced currents), the general form of the unsteady homogeneous heat conduction equation for an isotropic medium in cartesian coordinates reads [12]:
$\varrho c_{\text{p}}\frac{\partial T}{\partial t}=\nabla\cdot(\kappa\nabla T),$ (1)
with $\varrho$ being the mass density ($\text{kg}\cdot\text{m}^{-3}$), $c_{\text{p}}$ the specific heat capacity at constant pressure ($\text{J}\cdot\text{kg}^{-1}\cdot\text{K}^{-1}$), and $\kappa$ the heat conductivity ($\text{W}\cdot\text{m}^{-1}\cdot\text{K}^{-1}$).
Crucially, if $\kappa=\kappa(T)$ or it explicitly depends on space (e.g. multi-material PFC), then $\nabla\cdot(\kappa\nabla T)=\nabla\kappa\cdot\nabla T+\kappa\nabla^{2}T$, and a careful differentiation would be required. An elegant solution to this is employed in Ref. [49], which however forces an analytical form for the temperature-dependent thermal properties to be assumed. But for the specific case of ST40, the divertor heating phase during the plasma flat top lasts around 50 ms at most ($[90;140]\;\text{ms}$, Sec. 2.3). Therefore:
  • Assumption #1: the material properties are assumed to be temperature-independent.
  • Assumption #2: the divertor tiles are assumed to be a single layer of Mo of 29 mm in thickness, instead of the actual 4 mm of plasma-facing Mo overlaid to 25 mm of CuCrZr.
With respect to Asm. #2, the region affected by a temperature variation in a semi-infinite solid [^1] With $L=29$ mm of the present case, a reasonably constant heat flux over $[90;140]$ ms (Fig. 6), and assuming $\text{Fo}=\alpha t/L^{2}<0.2$ [12], the approximation of semi-infinite solid would hold for up to $\sim 3$ s. is here indeed limited to $2\sqrt{\alpha\cdot 50\;\text{ms}}\sim 3.3$ mm (with $\alpha\sim 54\;\text{mm}^{2}\cdot\text{s}^{-1}$ defined in Eq. (2)) along the depth of the tile, hence still confined within the only Mo stratum. No impurity surface layer [14, 49] is accounted for.
Asms. #1 and #2 imply model errors, which are formally assessed in Sec. 4.2.2. Eq. (1) accordingly reduces to:
$\frac{\partial T}{\partial t}=\alpha\nabla^{2}T,$ (2)
where $\alpha=\kappa/(\varrho c_{\text{p}})$ is the thermal diffusivity ($\text{m}^{2}\cdot\text{s}^{-1}$), unaffected by any differential operator only because of its independence on $T$ in the present case.
@@
The orthogonal reference frame adopted in FAHF is pictured in Fig. 1b and chosen such as to be aligned with the sides of the tiles. In particular, $x$ is the direction along the divertor (henceforth referred to as ”poloidal direction”), $y$ the direction along the depth of the tile, and $\phi$ is orthogonal to $x$ and $y$, and happens to be a proxy for the toroidal direction (which would actually be aligned to the strike point band). In ST40:
  1. 1.
    Assumption #3: the two-dimensional description of heat conduction is assumed to be appropriate, and the toroidal direction is neglected by assuming toroidal symmetry.
A bundle of $N_{\phi}$ independent parallel observation chords is generated, one such being shown in Fig. 1b (blue dots). The chords are aligned with the $x$ direction and sit at uniformly-spaced locations $\{\phi_{k}\}_{k=1}^{N_{\phi}}$ on the rectangular planar surface (Sec. 2.1) of the divertor tile CAD ($x\phi$ plane). This CAD is a suitable representation of the actual tiles, as long as surface imperfections are absent. Surface intrusions/protrusions deterministically cause noticeable abnormalities in the results. Any chords affected is therefore discarded from the analysis.
Each chord is then further divided along $x$ in two equal sub-chords, one per tile, where the gap running along $\phi$ lies. This is allowed by the thermal insulation provided by the gap (Sec. 3.1.4). The resulting 2D computational domain in FAHF is therefore a (32 mm) $\times$ (29 mm) rectangle in the $xy$ plane. The overall output along the entire chord (e.g. surface heat flux profile, Sec. 3.1.6) is ultimately obtained by juxtaposing in space the FAHF output of each of the two sub-chords at every time-step, as accomplished for $T_{\text{IR}}$ in Fig. 1c.
Asm. #3, relying on FAHF being independently run along (each sub-chord of) each chord, is justified by the fact that the poloidal temperature gradients $\partial_{x}T_{\text{IR}}$ dominate over the toroidal gradients $\partial_{\phi}T_{\text{IR}}$ when the strike point lies in the camera FoV, i.e. the time window of most interest $[90;140]\;\text{ms}$. However, following from the peculiar three-dimensionality of the ST40 divertor, $\partial_{\phi}T_{\text{IR}}\neq 0$ holds, as evident from Fig. 1b. Therefore, although FAHF’s multi-2D method is still preferred over a full 3D treatment for reasons of computational cost, the geometrical error implied by the multi-2D approach is assessed in Sec. 4.2.1.

### v2

@@ Abstract
Infra-red (IR) thermography is an essential diagnostic tool for understanding the edge plasma behavior in fusion devices. In this work, we present a new in-house numerical tool, Functional Analysis of Heat Flux (FAHF), for IR thermographic inversion on Tokamak Energy’s spherical tokamak (ST40). FAHF, written in Python, is designed for multi-2D thermographic inversions by solving the heat conduction equation within the divertor tiles using the finite difference method, and an explicit time stepping scheme. Utilising IR camera data with the highest available effective spatial resolution, FAHF calculates the plasma perpendicular heat flux density on the divertor tile surfaces -- a crucial quantity for edge plasma analysis. The tool’s internal numerics is first verified through formal time and space convergence analyses, and further corroborated by an energy balance assessment. Although FAHF demonstrates significant sensitivity to user-selected spatial resolution, precise heat flux values are recoverable by ensuring a sufficiently high resolution. Implications for the optimal resolution of both the code and the diagnostic system are discussed. Finally, FAHF’s model and geometry simplifications are confirmed to be accurate within 10%, based on comparison with COMSOL Multiphysics® simulations. As such, FAHF is proven to be a precise and accurate tool for IR thermographic inversions in ST40.
@@ Introduction
Tokamak-based nuclear fusion presents numerous physico-technological challenges. Among them, a robust power exhaust solution which maintains the integrity of plasma-facing components (PFCs) is critical[1]. Understanding the scrape-off layer (SOL) transport [2] is essential for establishing such a robust solution. In pilot plants [3, 4], unmitigated heat flux densities of hundreds of $\text{MW}\times\text{m}^{-2}$ are expected. Despite the mitigation strategies routinely undertaken in tokamaks [5, 6, 7], the current PFC technologies[8] are only effective up to tens of $\text{MW}\times\text{m}^{-2}$, therefore necessitating a reliable determination of heat flux parameters such as peak heat flux $q_{0}$ and power fall-off length $\lambda_{q}$ for appropriate designs.
Infra-red (IR) thermography[9] emerges as a prime diagnostic tool for characterising the SOL plasma and determining these heat flux parameters. The evolution in time of the PFC surface temperature, measured via the IR thermography, informs the IR thermographic inversion, which involves the solution of the heat conduction equation[10]. Although this is a well-established method, complexities such as space and temperature dependencies of PFC material properties, and their intrinsic three-dimensionality[11] complicate modelling efforts, also by lengthening the required computational time.
Numerous teams have significantly contributed to the development of IR thermographic systems and inversion techniques. Among them are: Alcator C-Mod[12], ASDEX[13, 14, 15], COMPASS[16], DIII-D[17, 18], EAST[19], ITER[20], JET[21, 22], KSTAR[23, 24], LHD[25], MAST[26, 27], NSTX[28, 18], SPARC[29, 30], TCV[31, 32], W7-X[33], and WEST[34, 35, 36].
Recent experiments, including those on the ST40 spherical tokamak[37], owned and operated by Tokamak Energy Ltd., COMPASS[38] and TCV[39] suggest that predictions from existing scalings[40] may overestimate $\lambda_{q}$ in certain scenarios, hence rising concern on the appropriate spatial resolution of the diagnostic system. An analysis of the above-mentioned references indicates that the hardware specifications at times only marginally meet, or fail to meet, the resolution requirements needed to accurately diagnose $\lambda_{q}$, even for values predicted by the existing scalings[40].
In this work, we introduce the high-resolution IR thermography system of ST40 and FAHF (Functional Analysis of Heat Flux), a new in-house Python-based flexible tool for high-resolution IR thermographic inversions on ST40. FAHF addresses the afore-mentioned modelling challenges with a fast, precise and accurate approach for solving the heat conduction equation. It is developed with prime emphasis on speed and modularity, which allows model extensions and applications to different cases.
FAHF’s precision is verified through formal convergence analyses. The accuracy of the tool is then estimated via COMSOL Multiphysics®, by means of a cross-code comparison, a proven method for quantifying model and geometry errors [41, 42], paramount in ST40. We also demonstrate that ST40’s IR thermography is particularly well-suited for diagnosing both a narrow[37] $\lambda_{q}$ and fine-scale structures in the heat flux footprint.
@@
The paper is organised as follows. Sec. 2 expounds on the ST40 tokamak, experimental setup and data selection. The fundamentals of the thermographic inversion are then described in Sec. 3. The numerical methods and tools adopted in this work are introduced in Sec. 4. The results of the study are collated in Sec. 5, and discussed in Sec. 6. Lastly, Sec. 7 summarises the conclusion of the work, and comments on future studies.
@@
This section aims at detailing the relevant features of the ST40 tokamak, of its divertor and of the IR thermographic apparatus. With the identification of a suitable plasma shot in the ST40 database, this ultimately constitute a solid foundation for a high-resolution thermographic inversion.
@@
ST40 [43, 44] is a compact, high-field spherical tokamak, with a major radius $0.4-0.5$ m, minor radius $0.2-0.25$ m, and a magnetic field up to 2.0 T on axis.
Disconnected double-null diverted plasmas [45, 46] are most commonly featured in ST40. Fig. 1a shows the EFIT magnetic reconstruction [47] in the poloidal cross section for shot 11376 at time instant 128 ms (Sec. 2.3). Its primary (solid) and secondary (dashed) separatrices are distinct, and displayed in magenta. In Fig. 1a, the secondary separatrix only contacts the wall (black) with its divertor legs above the X-point, leaving enough clearance to the centre column ($R=0.17$ m).
ST40 is provided with highly-shaped, non-axisymmetric, fish-scaled[11] tiled divertor targets[48], pictured in Fig. 1b, with inter-tile gaps of 4.5 mm in width, and no active cooling. The leading edge of each tile is poloidally chamfered[11], an area excluded in the present study. The leftover volume of the divertor tiles is thus a rectangular cuboid, with 4 mm of plasma-facing TZM molybdenum (Mo) brazed onto 25 mm of copper-cromium-zirconium (CuCrZr).
Poloidally, the geometry of the divertor is that of a short-legged, open, non-baffled divertor, with the inner leg orthogonal to the plate, and an horizontal outer leg (Fig. 1a). This configuration suppresses the role played by neutral particles at the outer target[49, 50, 51], hence favouring attached plasmas in ST40. This is required for IR thermographic inversions leading to meaningful $\lambda_{q}$ estimates[40]. Additionally, the ST40 divertor geometry allows a strong optimisation of the diagnostic system, instead intrinsically more challenging in presence of a closed divertor with vertical targets [12].
@@
Hardware characteristics
ST40 is equipped with a FLIR X6903sc MWIR ($3-5\;\mu$m) camera, with an acquisition rate of 0.8 kHz ($\Delta t_{\text{IR}}=$ 1.25 ms) for its 612 $\times$ 540 pixels, and diagnoses the upper outer divertor target (Fig. 1a). The reader is redirected to Refs.[9], [14] and [18] for the fundamental principles of IR thermography, and to Table 1 for the specifications of the ST40 system, further commented below.
Table 1: Specifications of ST40’s FLIR X6903sc MWIR camera: diagnosed wavelength spectrum; acquisition rate; time resolution; pixels along $\phi_{\text{p}}$ and $x_{\text{p}}$ directions (Fig. 1b); average spatial resolution along the divertor; field of view (FoV) along $\phi_{\text{p}}$ and $x_{\text{p}}$ in degrees (deg); angle between the line of sight (LoS) and divertor surface normal; calibrated dynamic temperature range; saturation temperature.
Spectrum ($\mu$m) Rate (kHz) $\Delta t_{\text{IR}}$ (ms) Pixels (--) $\Delta x_{\text{IR}}$ (mm/pix) FoV (deg) LoS angle (deg) Range (℃) Saturation (℃)
$3-5$ 0.8 1.25 640 $\times$ 512 0.21 $\pm$ 0.03 ($14.6;12.7$) 38 [$12;106$] 118
@@
Figure 1: Shot 11376 of the ST40 database at 128 ms. (a) EFIT reconstruction of the disconnected double-null magnetic equilibrium: primary and secondary separatrix in magenta solid and dashed, respectively; further magnetic surfaces in grey; wall in black; CH#2 in blue (1:1 scale); LoS in dashed blue. (b) IR camera view on the upper outer divertor in pixel space $(x_{\text{p}},\phi_{\text{p}})$. Overlaid are (see text for details): the FAHF reference frame in real space with $\phi$, proxy for toroidal direction, $x$, proxy for poloidal direction, and the direction in the depth of the tiles $y$ (entering the screen); an illustrative representation of the chords CH#1 and CH#2 (blue dots) of the bundle in both their sub-chords, one per tile. (c) Temperature $T_{\text{IR}}$ as extracted from the IR view via the Calcam software [52] (Appendix Calcam) along CH#2 in real space, with the empty gap in its middle. The data-point resolution is the actual $0.21\pm 0.03$ mm/pixel of the ST40 IR camera.
Figure 2: Main components of the endoscope assembly in ST40: camera protective housing (CAM); outer vacuum chamber (OVC); zoomed-in endoscope support (ES); schematic of the inner vacuum chamber (IVC), further detailed in Fig. 1a.
An ad-hoc re-entrant endoscope has been designed for ST40 with Zinc Selenide (ZnSe) optics, and is pictured in Fig. 2. The horizontal sub-assembly (top) contains the camera protective housing (CAM) and is fixed to the vertical sub-assembly, penetrating the outer vacuum chamber (OVC), by means of a connection flange. The assembly terminates with endoscope support (ES), facing the inner vacuum chamber (IVC).
The resulting camera line of sight (LoS, dashed blue line in Fig. 1a) within the IVC is at an acceptable angle [^1] Data related to LoS, FoV and pupil position are extracted via Calcam[52] (Appendix Calcam). of 38 degrees to the divertor surface normal (LoS angle). Instead, the field of view (FoV, 14.6 degrees along $\phi_{\text{p}}$ and 12.7 degrees along $x_{\text{p}}$ in Fig. 1b) is such that to cover the minimum amount of surface required to fully capture the heat flux footprint. The compenetration of the above LoS, FoV, short pupil-divertor distance (0.37 m) and the accommodating divertor geometry (Sec. 2.1) results in a noteworthy effective spatial resolution of the camera, averaging $\Delta x_{\text{IR}}=0.21\pm 0.03$ mm/pixel ($\sim 5$ pixel/mm) over two divertor tiles, and appreciable in Fig. 1c. The comparison with the list of references in Sec. 1 actually suggests the spatial resolution of ST40 IR camera to be the highest available worldwide for thermographic inversions of a full plasma heat flux footprint.
Worthwhile stressing is also the orientation of the camera LoS relative to the plasma magnetic topology: the LoS prevalently crosses the optically-thin far SOL, and is thus less contaminated by plasma absorption/emission in the IR spectrum.
Calibration and surface properties
From the original IR camera dynamic range $[10;90]$, factory-calibrated with its 50-mm lens, and the in-house characterisation of the endoscope transmittance yielding 0.85, the resulting calibrated range for this work is $[12;106]$. The calibration is still applicable in $(106;118]$, with an accuracy however diminishing as the IR camera saturation temperature of $118$ is approached.
The surface emissivity of the divertor tiles is assumed to be 1.0, constant with respect to the position in space and surface temperature. Despite the molybdenum divertor surface, no spurious reflection is observed to affect the view.
Neither the calibration factor/range nor the surface emissivity impact the numerical characterisation of FAHF, focus of the present study. Also, Appendix Experimental uncertainty suggests that only a small correction on the output should be expected from an uncertainty of these parameters. A dedicated assessment of such factors will be accomplished, being instead required for any physical investigation of ST40 edge plasma.
@@
Shot number 11376 within the ST40 database is selected for the analyses. According to database-wide analyses[37], this is a representative ST40 plasma shot from the edge plasma point of view, and features a single-exponential heat flux footprint (Sec. 5.1.3). The absence of edge-localised modes [53] (ELMs) evolving on short time-scales simplifies the numerical investigations (Sec. 5.2.1).
@@
The plasma presents a stable flat-top during the time window $[50;140]$ ms, when the plasma is NBI-heated with $\sim 0.8$ MW of power. Although the IR camera is marginally affected by saturation in $[130;140]$ ms in the region of interest (Sec. 4.3.1), this does not impact the numerical assessment of FAHF.
@@
The magnetic equilibrium of shot 11376 (Fig. 1a) is nominally a lower-biased disconnected double-null (Sec. 2.1), i.e. the primary separatrix contacts the lower divertor, thereupon rerouting most of the power[45, 46]. However, the low degree of disconnection ($\delta r_{\text{sep}}/\lambda_{q}\sim 1.5$ at the outer mid-plane [54]) allows the plasma to still deposit a satisfactory amount of power on the upper divertor (in view of the camera) during the actual divertor heating phase in $[90;140]\;\text{ms}$.
Fundamentals of IR thermographic inversion
Solving the heat equation
By safely discarding any volumetric heat source/sink [^2] For instance, adding the volumetric Ohmic heating caused by the $\sim 1$ A of current drew by a Langmuir probe [55] in ST40 does not cause any appreciable change in the output., the general form of the unsteady homogeneous heat conduction equation for an isotropic medium in cartesian coordinates is given by [10]:
$\varrho c_{\text{p}}\frac{\partial T}{\partial t}=\nabla\cdot(\kappa\nabla T)\>,$ (1)
with $\varrho$ being the mass density ($\text{kg}\times\text{m}^{-3}$), and $c_{\text{p}}$ the specific heat capacity at constant pressure ($\text{J}\times\text{kg}^{-1}\times\text{K}^{-1}$). Lastly, $\kappa$ is the heat conductivity ($\text{W}\times\text{m}^{-1}\times\text{K}^{-1}$), which is acted upon by the divergence operator.
Upon imposing boundary conditions (BCs) in space, and an initial condition (IC) in time at $t=0$ ms, Eq. (1) is solved for the temperature $T$. According to the methodology specific to the thermographic inversion, a Dirichlet-type BC is enforced via the experimental temperature data $T_{\text{IR}}$ (Fig. 1c) at the plasma-facing boundary of the modelled domain. On the remaining non-plasma-facing boundaries, homogeneous Neumann BCs are assumed, physically representing adiabatic (no-flux) boundaries. This is legitimate in ST40 because of a negligible black-body emission in the gap ($T_{\text{IR}}<50$), and of a negligible heat convection/conduction provided by the rarefied neutral gas filling this volume (few Pa in pressure at most [56]). Additionally, the thermal contact of the tile back-surface is neglected. Applying various Dirichlet BCs at the back-surface is investigated, but does not cause any noticeable corrections in the time window $[90;140]$ ms, consistently with Sec. 4.1.1 and 6.4.1.
Surface heat flux and Eich fit
The temperature distribution $T$ from Sec. 3.1 informs the computation of the quantity of primary physical interest of the IR thermographic inversion, the heat flux vector at the PFC surface $(\text{W}\times\text{m}^{-2})$:
$\mathbf{q}=-\kappa\nabla T\>,$ (2)
and, specifically, its perpendicular component $\mathbf{q}\cdot\mathbf{n}$ (herein referred to as ”surface heat flux”) along the tile surface normal direction $\mathbf{n}$ (defined as entering in the tile). Once this quantity is calculated, the thermographic inversion is nominally concluded.
However, further information about the SOL plasma is customarily retrieved by fitting the surface heat flux profile in space with the Eich function $q_{\text{Eich}}$ from Eq. (1) of Ref.[40]. The main output of such fitting procedure is embodied by the parameters of: maximum heat flux $q_{0}$, power fall-off length $\lambda_{q}$, spreading factor $S$, background heat flux $q_{\text{BG}}$, strike point location $x_{0}$, and poloidal flux expansion $f_{X}$. Ref.[40] is recommended for detailed descriptions of such quantities.
The integral value of the Eich profile $q_{\text{Eich}}(x)$ along the divertor, proportional to the total power, is here additionally added, and defined as:
[…]
Although the formal iterative procedure of Ref.[48] is actually best-posed when fitting with the multi-parameter non-linear function $q_{\text{Eich}}$, and should be more widely adopted in the community, this is not here accomplished because of its computational cost. Instead, the fit is deemed satisfactory upon showing $R^{2}>98.5\%$, and is herein merely considered as a parametrisation of the surface heat flux, to extract insightful parameters to characterise FAHF’s precision. This simplified approach is therefore acceptable, and is justified a posteriori by Secs. 5.1.3 and 5.1.4, which show a genuine behaviour of the heat flux.
The interested reader is redirected to Refs. [48] and [37] for the actual quantification of $\lambda_{q}$ in ST40, which is outside the scope of the present work.
Internal energy variation and deposited energy
From Secs. 3.1 and 3.2, the internal energy variation over a time interval $\tau$ throughout the volume $V$ of the domain:
[…]
and the net energy over $\tau$ flowing through the tile surface $\mathbf{S}=S\times\mathbf{n}$:
[…]
are computed. This aids the assessment of the energy-preserving properties of the algorithms. With $\mathbf{n}$ defined in Sec. 3.2, $\mathcal{E}_{\text{surf}}>0$ when deposited in the tile.
@@
The present section provides essential details of the assumptions, choices and implementation of FAHF algorithms.
Model assumptions
Model assumptions are adopted in FAHF to simplify and speed up the solution of Eq. (1). Crucially, if $\kappa=\kappa(T)$ or it explicitly depends on space (e.g. multi-material PFCs), then $\nabla\cdot(\kappa\nabla T)=\nabla\kappa\cdot\nabla T+\kappa\nabla^{2}T$, and a careful differentiation would be required. An elegant solution to this is employed in Ref.[57], however not needed in the specific case of ST40, where the divertor heating phase lasts $\sim 50$ ms at most ($[90;140]\;\text{ms}$, Sec. 2.3) -- not enough to result in a temperature which would appreciably change the material properties. Therefore:
  • Asm. #1. The material properties are assumed to be temperature-independent.
No impurity surface layer [18, 58] is accounted for. Moreover, the region affected by a temperature variation in a semi-infinite solid [^3] With $L=29$ mm of the present case, a reasonably constant heat flux over $[90;140]$ ms (Fig. 6), and assuming $\text{Fo}=\alpha t/L^{2}<0.2$ [10], the approximation of semi-infinite solid would hold for up to $\sim 3$ s. is here limited to $2\sqrt{\alpha\times 50\;\text{ms}}\sim 3.3$ mm (with $\alpha\sim 54\;\text{mm}^{2}\times\text{s}^{-1}$ defined in Eq. (6)) along the depth of the tile, hence still confined within the only Mo stratum. Consequently:
  • Asm. #2. The divertor tiles are assumed to be a single layer of Mo of 29 mm in thickness, instead of the actual 4 mm of plasma-facing Mo overlaid to 25 mm of CuCrZr.
Asms. #1 and #2 imply minor model errors in FAHF, as demonstrated in Sec. 5.2.3, and Eq. (1) accordingly reduces to:
$\frac{\partial T}{\partial t}=\alpha\nabla^{2}T\>,$ (6)
where $\alpha=\kappa/(\varrho c_{\text{p}})$ is the thermal diffusivity ($\text{m}^{2}\times\text{s}^{-1}$), unaffected by any differential operator only by virtue of being assumed constant in the present case.
@@
The orthogonal reference frame used in FAHF is depicted in Fig. 1b and chosen as aligned with the sides of the cuboidal tiles. In particular, $x$ is the direction along the divertor (henceforth referred to as ”poloidal direction”), $y$ the direction along the depth of the tile, and $\phi$ happens to be a proxy for the toroidal direction (which would actually be aligned to the strike point band). When solving Eq. (6) in FAHF:
  • Asm. #3. The two-dimensional description of heat conduction is assumed to be appropriate, and the toroidal direction is neglected by assuming toroidal symmetry.
A bundle of $N_{\phi}$ independent parallel observation chords is generated, two such being shown in Fig. 1b (blue dots), with a more complete representation found in Ref.[48]. The chords are aligned with the $x$ direction and sit at uniformly-spaced locations $\{\phi_{k}\}_{k=1}^{N_{\phi}}$ on the rectangular planar surface (Sec. 2.1) of the divertor tile CAD ($x\phi$ plane). This CAD is a suitable representation of the actual tiles, as long as surface imperfections are absent. Surface intrusions/protrusions deterministically cause noticeable abnormalities in the results[48]. Any chords affected is therefore discarded.
Each chord is then further divided along $x$ in two equal sub-chords, one per tile, where the toridal-running gap lies. This is allowed by the thermal insulation provided by the gap (Sec. 4.1.4). The resulting 2D computational domain in FAHF is therefore a (32 mm) $\times$ (29 mm) rectangle in the $xy$ plane. The overall output along the entire chord (e.g. surface heat flux profile, Sec. 4.1.6) is ultimately obtained by juxtaposing in space the FAHF output of each of the two sub-chords at every time-step, as accomplished for $T_{\text{IR}}$ in Fig. 1c.
Asm. #3, resulting in FAHF being independently run along (each sub-chord of) each chord in a ”multi-2D” fashion, is expected to be justified by the fact that the poloidal temperature gradients $\partial_{x}T_{\text{IR}}$ tends to dominate over the toroidal gradients $\partial_{\phi}T_{\text{IR}}$ in the majority of plasma properties. However, following from the peculiar three-dimensionality of the ST40 divertor, $\partial_{\phi}T_{\text{IR}}\neq 0$ still holds, as evident from Fig. 1b, and also detailed in Ref.[48]. The geometrical error implied by the multi-2D approach, instead of a 3D one, is assessed in Sec. 5.2.1, and shown to be within an acceptable limits in the bulk of the tile.
Compared to the ”single-chord” approach most commonly employed within the community, this multi-2D approach also beneficially provides statistics and relevant error-bars in the value of $\lambda_{q}$ ultimately estimated [48]. Moreover, the multi-2D treatment properly accounts for difference of the perpendicular heat flux on a chord-to-chord basis, crucial for a power balance analysis, particularly in ST40.